F(x) = (3sinx – 4cosx) – 7 how to find its maximum and minimum value

Grade:12th pass

1 Answers

Arun
6 years ago
we should know that 
for anyepression like 
a sinx + b cosx  OR  a sinx –b cosx  
maximum value is = \sqrt{}(a2 + b2)
while minimu value is =  \sqrt{}(a2 + b2)
hence for the given question
maximum value is = \sqrt{}(32 + 42) – 7 = 5 –7 = –2
minimum value is = – \sqrt{}(32 + 42) – 7  =  –5 – 7 = –12